OpenStudy (kash_thesmartguy):

Am I doing this correctly: (2^3)^4 2^12 ??? Please help.

OpenStudy (candyme):

\[(2^{3})^{4} = 3^{(3 \times 4)}\]

OpenStudy (candyme):

I'm sorry I wrote 3 for some reason. It should be \[2^{12} = 4096\]

OpenStudy (anonymous):

yah its 4096

OpenStudy (unklerhaukus):

\[(2^3)^4\\ =(2\times2\times2)\times(2\times2\times2)\times(2\times2\times2)\times(2\times2\times2)\\ =2\times2\times2\times2\times2\times2\times2\times2\times2\times2\times2\times2\\ =2^{12}\]

OpenStudy (anonymous):

in other word you are doing it correctly